### v1

@@ Abstract
Localized surface plasmons can confine light within a deep-subwavelength volume comparable to the scale of atoms and molecules, enabling ultrasensitive responses to near-field variations. On the other hand, this extreme localization also inevitably amplifies the unwanted “noise” from the response of local morphological imperfections, leading to complex spectral variations and reduced consistency across the plasmonic nanostructures. Seeking uniform optical responses has therefore long been a sought-after goal in nanoplasmonics. However, conventional probing techniques by dark-field (DF) confocal microscopy, such as image analysis or spectral measurements, can be inaccurate and time-consuming, respectively. Here, we introduce SPARX, a deep-learning-powered paradigm that surpasses conventional imaging and spectroscopic capabilities. In particular, SPARX can batch-predict broadband DF spectra (e.g., 500–1000 nm) of numerous nanoparticles simultaneously from an information-limited RGB image (i.e., below 700 nm). It achieves this extrapolative inference beyond the camera’s capture capabilities by learning the underlying physical relationships among multiple orders of optical resonances. The spectral predictions only take milliseconds, achieving a speedup of three to four orders of magnitude compared to traditional spectral acquisition, which may take from hours to days. As a proof-of-principle demonstration for screening identical resonances, the selection accuracy achieved by SPARX is comparable to that of conventional spectroscopy techniques. This breakthrough paves the way for consistent plasmonic applications and next-generation microscopies.
@@
Deep learning, Dark-field, Near-field, Far-field, Single plasmonic nanocavity, Consistency, Predicting spectra from images
@@ Introduction
Light weakly interacts with microscopic entities such as atoms and molecules due to their vast dimensional mismatch (e.g., wavelength $\simeq 10^{2-3}$ nm versus size $\simeq 10^{1}$). Plasmonic nanogaps—formed by closely spaced metallic surfaces—support localized surface plasmons that confine light to the deep-subwavelength scale, bridging this mismatch [1, 2, 3, 4, 5]. By integrating low-dimensional materials into these nanogaps, plasmonic systems enable highly sensitive detection of electronic (e.g., photoluminescence) [6], vibrational (e.g., surface-enhanced Raman scattering) [7], and other linear and nonlinear optical properties of materials [8, 9]. This high sensitivity is usually attributed to the extremely strong and localized field that enhances the light-matter interaction. [10]
@@
However, this localization also amplifies the information from the local morphologies such as nonuniformity, roughness, defects, surface atomic-level dynamics and pico-cavities, inevitably introducing noise and compromising reproducibility [11, 12, 13, 14, 15, 16, 4, 17, 18]. Even monocrystalline nanoparticles with solution-phase spectral uniformity could exhibit unpredictable resonance broadening when interfaced with metallic films to form plasmonic nanogaps [19]. This intrinsic variability creates a paradox: maximizing field enhancement may inherently compromise the spectral consistency of the plasmonic systems. One way of circumventing this is to screen the nanoparticles with identical resonances from countless particles ($\simeq 10^{10}$ per milliliter in the solution) drop-casted at once on the substrate during the sample preparation, and only perform experiments on those spectrally-uniform structures.
Confocal dark-field (DF) microscopy has traditionally served as a primary detection technique in plasmonics and biosensing, enabling simultaneous spatial and spectral characterization at the single-nanoparticle level. Researchers typically assess the uniformity of the sample and infer the resonance of plasmonic systems based on an empirical assumption—that the color distribution of Airy patterns in DF images has direct correlation with the energy and the lineshape of the resonances. However, human vision may not resolve subtle chromatic variations, and the displayed color based on RGB multiplexed channels of commercial color cameras (400-700 nm) often lacks sufficient spectral information to capture lower-energy resonances. Moreover, even within the detection band, RGB encoding can significantly reduce spectral precision by compressing information into three digits. Furthermore, the presence of multiple resonances with different energies and far-field radiation patterns [20, 21] adds extra complexity to the Airy patterns, making naive inference impossible. Therefore, although intrinsic correlations do exist between resonances across frequency bands and between DF images and spectra, the underlying relationships are often highly nonlinear and complex, posing challenges for conventional analytical approaches. This highlights the need for a simple, fast, and intelligent approach that can learn the correlations and accurately extrapolate beyond the camera’s physical limit.
@@
In this regard, deep learning (DL) has emerged as a transformative tool for decoding complex photonic interactions [22]. It has demonstrated the ability to learn intricate light–matter correlations through inverse design of plasmonic devices and metasurfaces [23, 24, 25, 26, 27, 28, 29]. When integrated with microscopic and nanoscopic techniques, DL has also shown success in noise suppression, signal processing, nanoparticle identification, and super-resolution imaging, which resolves variations and uncertainties at the subwavelength scale [30, 31, 32, 33, 34, 35, 36, 37, 38]. Overall, DL is particularly valuable for managing uncertainties in experimental and fabrication processes—such as material inconsistencies, structural imperfections, and environmental fluctuations. By leveraging large datasets, DL models enhance robustness and generalization, leading to more reliable analysis and design in plasmonic systems.
@@
Here, we present a multimodal DL framework, named SPARX (Spectral Prediction and Reconstruction from RGB with eXtrapolation), that captures the deeper correlations between DF images and broadband spectral responses, overcoming the inherent imprecision of human visual perception, the limited information in RGB channels, and the speed constraints of conventional optical characterization. We conducted automated measurements on more than 12,000 gold nanoparticles-on-mirror (NPoM) nanostructures, compiling a dataset that can capture the complexity of nanogap plasmons. In particular, most spectral resonances occur beyond 800 nm, demonstrating that the SPARX model can infer lower-order resonances from higher-order features, effectively extending spectral predictions beyond the physical capture capabilities of the camera. This can hardly be achieved with simple correlations or empirical experiences, especially with the same high level of precision. Furthermore, by leveraging heteroscedastic loss, the SPARX model could estimate the uncertainty in spectral predictions for individual nanoparticles, effectively quantifying the prediction error. This correlation allows for further refinement in the selection process. Our approach achieves a four-order-of-magnitude acceleration in characterization speed compared with spectral acquisition, enabling real-time classification of nanostructures in the future. By replacing spectrometer-dependent workflows with camera-based deep learning, we establish a scalable platform for next-generation nanophotonic device engineering, biosensor development, and quantum optical system design, where both spectral precision and high-throughput characterization have proven indispensable.
@@ Results
Among all plasmonic systems, the nanoparticle-on-mirror (NPoM) configuration (Figure 1(a)) has been widely observed as one of the most versatile platforms, offering easily controllable nanogaps for enhancing optical phenomena [3]. In our experiment, the NPoM system was formed by drop-casting 80 nm cetyltrimethylammonium (CTAC)-capped gold nanoparticles onto a gold mirror (see Methods for sample preparation). The nanogap was defined by a 1-2 nm thick CTAC molecule layer. From electromagnetic (EM) theory, it is well established that the size and shape of the nanoparticles [13, 17], small variations in the roughness of the metallic surface [12, 11], and the morphology of the gap between the nanoparticles and the metallic substrate (e.g., facets size and adatoms) [18, 15] can drastically alter the resonance behavior of the system.
@@
The attempt to uncover the correlations among different orders of the plasmonic resonances can be dated back to the Mie theory [39], where an analytical solution of the scattering from a spherical nanoparticle can be derived by solving the Maxwell equations. Each Mie resonance (i.e., dipole, quadrupole, etc.) can be intercorrelated through the spherical harmonics. When coupled with its own image through a mirror, the NPoM plasmonic nanogap can give rise to a richer variety of the hybridized modes, with the resonance of $mn$-order mode $\lambda_{mn}$ given by the cylindrical Bessel functions [18]: $\lambda_{mn}=\frac{\pi wn_{\mathrm{eff}}}{J_{mn}-\phi}$, where $J_{mn}$ is $n$-th root of the $m$-th order Bessel function. $n_{\rm eff}$ is the effective refractive index of the metal-insulator-metal junction, and $\phi$ is a proper reflection phase. However, this simplified model can only predict the cavity-like resonances of a perfectly spherical nanoparticle with a single ideal circular facet at the bottom gap, with uncertainties such as polyhedral shapes, roughness, nonlocal and quantum effects, and others mentioned above falling outside its scope. Thus, using a simplified, classical analytical solution to extend the spectra of a realistic system beyond the detection limit can be challenging.
@@
Figure 2: Unsupervised learning of the dark-field information. (a) Optical setup for DF images and spectra collection. WL: white light source. BS: beam-splitter with reflection : transmission in % (R:T)= 50:50, OBJ: Olympus objective, numerical aperture = 0.9, working distance = 1mm. (b) PCA explained variance, (c, d) First four principal components for DF images and spectra. (e) 1D and 2D PCA projections showing correlations between DF and spectral data, with DF projections colored by the index of the samples (timeline), PCA of spectra, peak location ($\lambda$), and integrated energy. (f) Similar analysis with 1D and 2D UMAP projection.
@@
To further support these observations, we applied Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality reduction technique, to both spectral and DF image data, as shown in Figure 2(f). The spectral data is embedded into a one-dimensional UMAP space, while the DF images are projected into two dimensions. By coloring the DF UMAP projection using various spectral-derived features—including UMAP values of the spectra, peak positions, and total spectral energy—we uncover strong visual correlations between the DF and spectral modalities (2D maps in Figure 2(f)). An interesting outcome is the appearance of multiple distinct clusters in the DF UMAP space (in contrast to PCA projection in Figure 2 (e)), which exhibit clear correspondence to specific spectral features. For example, the cluster located at the bottom right of the projection space shows a color gradient where resonance wavelength and total spectral energy vary in opposite directions—i.e., as the resonance red-shifts, the intensity decreases. Additionally, when coloring the projections based on sample index—correlated with the data acquisition time—we observe no discernible pattern, confirming the high reproducibility and stability of the measurements across multiple days. This improved clustering allows us to reveal the rich spectral information that DF images inherently carry, which is typically inaccessible In traditional DF imaging analysis.
@@
In contrast, the intensity distribution, shown in the right panel of Figure 3(b), presents a distinctly different behavior. The overall population density plot of resonance intensity (shown in black) displays three distinct peaks, indicating the existence of three subpopulations. Interestingly, the core group is primarily composed of the subpopulation with the lowest resonance intensities. The remaining two subpopulations with higher intensities are predominantly classified as outliers. Interestingly, the spectra with modest intensity ($\simeq$2.6 arb. units), which contain the largest number of samples, are most recognized as the core data that SPARX is most confident in. This may be attributed to their generally more uniform and well-behaved geometries, with reduced randomness stemming from previously discussed factors. In contrast, the higher-intensity subpopulations are treated as outliers by SPARX, likely due to more complex features such as sharp polyhedral facets, nonuniform gaps, or even picocavities. These features have more randomness in general, and are more difficult to characterize using only DF images, leading to increased uncertainty in their prediction. In fact, this also aligns with the empirical intuition in SERS: extreme enhancement with high intensity often compromises robustness and consistency. Nanoparticles exhibiting such exceptionally high intensities remain rare outliers relative to the overall population. Remarkably, our SPARX model captures this trade-off by analyzing the associated uncertainty. It is important to note that while our heteroscedastic model is capable of identifying such complex cases from DF images, accurately reconstructing their spectral details remains challenging due to the limitations of the input modality. Detection of these outliers should not be conflated with precise spectral prediction.
@@
In addition, Figure 4 showcases 23 examples (which are the same ones predicted by the homoscedastic model as shown in the SI, Fig. S3 from the test dataset), each accompanied by its corresponding DF image, SPARX heteroscedastic-predicted spectrum, and ground truth spectrum. The $y$-axis (intensity) of each DF spectrum is fixed to the same range, illustrating that SPARX can faithfully reconstruct both spectral intensity and wavelength with minimal errors. The predicted uncertainty is shown as blue shaded bands around the mean prediction, representing the standard deviation. Notably, when the predicted uncertainty is considered, the ground-truth spectra largely fall within this range. It means that, for the same DF images, the heteroscedastic SPARX model can resolve single-particle-level uncertainties.
@@
The theoretical upper limit of nanoparticles processable per FOV in our setup is estimated based on the non-overlapping spectral collection area of $\simeq$ 1.2 $\mu$m2 per nanoparticle (3–4 pixels of the Horiba spectrometer, with one pixel calibrated as $\simeq$ 0.3 $\mu$m). Under optimal conditions in Figure 5(c), $\sim$1000 nanoparticles can be processed in a single DL prediction cycle. Processing $\sim$1000 DF images with GPU takes only  0.4 s, suggesting a potential four-order magnitude speed improvement compared to conventional spectroscopy.

### v2

@@ Abstract
Localized surface plasmons confine light within deep-subwavelength volumes, enabling ultrasensitive near-field responses that underpin a wide spectrum of interdisciplinary technologies. Yet this extreme localization also amplifies unwanted “noise” from local nanomorphological variations, resulting in spectral complexities and inconsistencies that have long hindered reproducible and scalable nanophotonics. In this context, optical identification and screening of nanostructures with consistent, target responses offers a practical strategy. However, conventional imaging and spectroscopies, including hyperspectral methods, are limited by a resolution–throughput trade-off, motivating the development of faster, high-precision approaches. Here, we introduce SPARX, a deep-learning (DL)-powered paradigm that surpasses conventional imaging and spectroscopic capabilities. SPARX batch-classifies the nanoparticles by their shapes, and extrapolates broadband dark-field spectra (500–1000 nm) of numerous nanoparticles simultaneously from an information-limited RGB image (<<700 nm) by learning physical relationships among multiple orders of resonances. Predictions take only milliseconds, achieving a speedup of 2–4 orders of magnitude over traditional methods, while maintaining comparable precision. This transformative, imaging-DL integrated approach enables reproducible nanoplasmonic applications and, more importantly, fundamentally reshapes optical characterization workflows and extends their reach.
@@
Deep learning, Dark-field, Near-field, Far-field, Single plasmonic nanocavity, Consistency, Predicting spectra from images, Classification
@@ Introduction
Plasmonic nanostructures, particularly those forming nanogaps, support localized surface plasmons that confine light to deep-subwavelength volumes comparable to molecular and atomic dimensions. [1, 2, 3, 4, 5]. By integrating low-dimensional materials into these nanogaps, they enable probing of electronic (e.g., photoluminescence) [6], vibrational (e.g., surface-enhanced Raman scattering) [7, 4], and other linear and nonlinear optical properties of materials [8, 9], with atom-level sensitivity [10], introducing a fascinating field as ‘extreme nanophotonics[3, 11]. This high sensitivity arises from the extremely enhanced nanoscopic light-matter interactions supported by nanostructure’s quasinormal modes [12], making resonance–material spectral matching a fundamental prerequisite for nanophotonic applications.
@@
However, this extreme localization also amplifies contributions from local nanomorphological features—such as nonuniformity, roughness, defects, surface atomic-level dynamics, and pico-cavities—inevitably introducing ‘noise’ and compromising reproducibility [13, 14, 15, 16, 12, 17, 4, 18, 19]. This intrinsic variability [20, 3] poses a paradox—enhancing fields for exceptional single-particle performance can inherently undermine spectral consistency and reproducibility across ensembles, constituting a crucial bottleneck in extreme nanophotonics. One general way of circumventing this is to identify and pre-screen the target (e.g., spectrally-uniform, with desired resonances) nanoparticles from countless particles ($\simeq 10^{10}$ per milliliter in the solution) drop-casted on the substrate during the sample preparation, and conduct experiments only on that subset. Conventionally, this screening process has relied on manual expertise combined with imaging and spectroscopic measurements, rendering it inefficient and labor-intensive.
Confocal dark-field (DF) microscopy has long been a cornerstone for single-nanoparticle characterization. Resonances and sample uniformity are often inferred empirically from their images at high throughput, assuming that the color distribution of Airy patterns directly reflects the energy and lineshape of the plasmonic modes. However, human vision cannot reliably resolve subtle chromatic variations, and the information-compressed three-digit RGB output of commercial cameras (400–700 nm) lacks sufficient spectral detail and fails to capture lower-energy resonances. Multiple resonances with distinct far-field patterns further complicate the problem [21, 22]. More precise techniques, such as DF mapping, acquire spatially resolved spectra via point-scanning, representing a form of hyperspectral imaging (HSI) [23, 24, 25, 26, 27, 28]. Such sequential acquisition is unsuitable for large-field detection, which can be mitigated with alternative HSI strategies, including line-scanning (push-broom), wavelength-scanning [23], and snapshot techniques [28]. Recent advances, including compressive sensing [29, 30, 31, 32], have further improved acquisition speed. However, HSI methods inherently face a resolution–throughput trade-off: higher spectral resolution requires narrower channels, reducing photon flux per channel [33, 27]. This lowers the signal-to-noise ratio and, unless compensated by stronger illumination or more sensitive detectors, demands longer exposure times—especially critical for weak signals such as single nanoplasmonic structures. They also entail costly and specialized hardware, large data storage and copious processing requirements, and rigorous calibration protocols [24, 34]. In contrast, confocal DF microscopy combined with RGB imaging remains economical and effective, while readily compatible with emerging techniques (e.g., machine learning, compressive sensing). However, to fully leverage this approach for high-throughput, precise acquisition of plasmonic nanostructures, the complex and nonlinear correlations between DF images and spectra—particularly beyond the RGB camera’s physical limits—must be revealed.
@@
In this regard, deep learning (DL) has emerged as a transformative tool for decoding complex photonic interactions [35]. It has demonstrated the ability to learn intricate light–matter correlations through inverse design of plasmonic devices and metasurfaces [36, 37, 38, 39, 40, 41, 42, 43]. When integrated with microscopic and nanoscopic techniques, DL has also shown success in noise suppression, signal processing, nanoparticle identification, and super-resolution imaging, which resolves variations and uncertainties at the subwavelength scale [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Overall, DL is particularly valuable for managing uncertainties in experimental and fabrication processes—such as material inconsistencies, structural imperfections, and environmental fluctuations.
@@
Here, we present a DL framework, named SPARX (Spectral Prediction and Reconstruction from RGB with eXtrapolation), that captures the deeper correlations between information-limited DF images and broadband spectral responses. It overcomes the inherent imprecision of human visual perception, the limited information in RGB channels, and the speed constraints of conventional optical characterization. SPARX leverages RGB imaging’s maximal photon efficiency (3 broad channels) to capture fast snapshots, and reconstructs the broadband spectra with high throughput (0.4 s/1,000 particles on GPU). In particular, most of the spectral resonances in our examples occur beyond 800 nm, demonstrating that the SPARX model can infer lower-order resonances from higher-order features by understanding their physical relationships, effectively extending spectral predictions beyond the physical capture capabilities of the camera. This can hardly be achieved with simple correlations or empirical experiences, especially with the same high level of precision.
Furthermore, by leveraging heteroscedastic loss, the SPARX model could estimate the uncertainty in spectral predictions for individual nanoparticles, effectively quantifying the prediction error. This correlation allows for further refinement in the selection process. Moreover, as a powerful complementary screening approach, we develop a SPARX classifier capable of directly identifying and differentiating nanoparticle shapes from DF RGB images—showcasing an inference capability beyond the reach of conventional spectral acquisition methods. All in all, our approach achieves a $10^{2}$$10^{4}$-fold acceleration in characterization speed compared with spectral acquisition, enabling real-time classification of nanostructures in the future. By replacing spectrometer-dependent workflows with camera-based deep learning, we establish a scalable platform for next-generation optical characterization, nanophotonic device engineering, and biosensor development, where both spectral precision and high-throughput characterization have proven indispensable.
@@ Results
Among all plasmonic systems, the nanoparticle-on-mirror (NPoM) configuration (Figure 1(a)) has been widely observed as one of the most versatile platforms, offering easily controllable nanogaps for enhancing optical phenomena [3]. Therefore, we select NPoM as a representative platform for demonstration, measuring 12,000 gold NPoMs to compile a dataset that captures the underlying complexity. In our experiment, the NPoM system was formed by drop-casting 80 nm cetyltrimethylammonium (CTAC)-capped gold nanoparticles onto a gold mirror (see Methods for sample preparation). The nanogap was defined by a 1-2 nm thick CTAC molecule layer. From electromagnetic theory, it is well established that the size and shape of the nanoparticles [15, 18], small variations in the roughness of the metallic surface [14, 13], and the morphology of the gap between the nanoparticles and the metallic substrate (e.g., facets size and adatoms) [19, 12] can drastically alter the resonance behavior of the system.
@@
The attempt to uncover the correlations among different orders of the plasmonic resonances can be dated back to the Mie theory [55], where an analytical solution of the scattering from a spherical nanoparticle can be derived by solving the Maxwell equations. Each Mie resonance (i.e., dipole, quadrupole, etc.) can be intercorrelated through the spherical harmonics. When coupled with its own image through a mirror, the NPoM plasmonic nanogap can give rise to a richer variety of the hybridized modes [56], with the resonance of $mn$-order mode $\lambda_{mn}$ given by the cylindrical Bessel functions [19]: $\lambda_{mn}=\frac{\pi wn_{\mathrm{eff}}}{J_{mn}-\phi}$, where $J_{mn}$ is $n$-th root of the $m$-th order Bessel function. $n_{\rm eff}$ is the effective refractive index of the metal-insulator-metal junction, and $\phi$ is a proper reflection phase. However, this simplified model can only predict the cavity-like resonances of a perfectly spherical nanoparticle with a single ideal circular facet at the bottom gap, with uncertainties such as polyhedral shapes, roughness, nonlocal and quantum effects, and others mentioned above falling outside its scope. Thus, using a simplified, classical analytical solution to extend the spectra of a realistic system beyond the detection limit can be challenging.
@@
Figure 2: Unsupervised learning of the dark-field information. (a) Optical setup for DF images and spectra collection. WL: white light source. BS: beam-splitter with reflection : transmission in % (R:T)= 50:50, OBJ: Olympus objective, numerical aperture = 0.9, working distance = 1mm. (b) PCA explained variance, (c, d) First four principal components for DF images and spectra. (e) 1D and 2D PCA projections showing correlations between DF and spectral data, with DF projections colored by the index of the samples (timeline), PCA of spectra, peak location ($\lambda$), and integrated energy.
@@
To further support these observations, we applied Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality reduction technique, to both spectral and DF image data, as shown in SI Figure S2. The spectral data is embedded into a one-dimensional UMAP space, while the DF images are projected into two dimensions. By coloring the DF UMAP projection using various spectral-derived features—including UMAP values of the spectra, peak positions, and total spectral energy—we uncover strong visual correlations between the DF and spectral modalities (2D maps in SI Figure S2). An interesting outcome is the appearance of multiple distinct clusters in the DF UMAP space (in contrast to PCA projection in Figure 2 (e)), which exhibit clear correspondence to specific spectral features. For example, the cluster located at the bottom right of the projection space shows a color gradient where resonance wavelength and total spectral energy vary in opposite directions—i.e., as the resonance red-shifts, the intensity decreases. Additionally, when coloring the projections based on sample index—correlated with the data acquisition time—we observe no discernible pattern, confirming the high reproducibility and stability of the measurements across multiple days. This improved clustering allows us to reveal the rich spectral information that DF images inherently carry, which is typically inaccessible in traditional DF imaging analysis.
@@
In contrast, the intensity distribution, shown in the right panel of Figure 3(b), presents a distinctly different behavior. The overall population density plot of resonance intensity (shown in black) displays three distinct peaks, indicating the existence of three subpopulations. Interestingly, the core group is primarily composed of the subpopulation with the lowest resonance intensities. The remaining two subpopulations with higher intensities are predominantly classified as outliers. Interestingly, the spectra with modest intensity ($\simeq$2.6 arb. units), which contain the largest number of samples, are most recognized as the core data that SPARX is most confident in. This may be attributed to their generally more uniform and well-behaved geometries, with reduced randomness stemming from previously discussed factors. In contrast, the higher-intensity subpopulations are treated as outliers by SPARX, likely due to more complex features such as sharp polyhedral facets, nonuniform gaps, or even picocavities. These features have more randomness in general, and are more difficult to characterize using only DF images, leading to increased uncertainty in their prediction. In fact, this also aligns with the empirical intuition in plasmonics and surface-enhanced Raman scattering: extreme enhancement with high intensity often compromises robustness and consistency. Nanoparticles exhibiting such exceptionally high intensities remain rare outliers relative to the overall population. Remarkably, our SPARX model captures this trade-off by analyzing the associated uncertainty. It is important to note that while our heteroscedastic model is capable of identifying such complex cases from DF images, accurately reconstructing their spectral details remains challenging due to the limitations of the input modality. Detection of these outliers should not be conflated with precise spectral prediction.
@@
In addition, Figure 4 showcases 23 randomly selected examples (which are the same ones predicted by the homoscedastic model as shown in the SI, Fig. S4 from the test dataset), each accompanied by its corresponding DF image, SPARX heteroscedastic-predicted spectrum, and ground truth spectrum. The $y$-axis (intensity) of each DF spectrum is fixed to the same range, illustrating that SPARX can faithfully reconstruct both spectral intensity and wavelength with minimal errors. The predicted uncertainty is shown as blue shaded bands around the mean prediction, representing the standard deviation. Notably, when the predicted uncertainty is considered, the ground-truth spectra largely fall within this range. It means that, for the same DF images, the heteroscedastic SPARX model can resolve single-particle-level uncertainties.
@@
Beyond NPoM, we further validated the generality of SPARX on a more complex plasmonic system—nanopatch antennas, composed of single $\sim$75 nm gold nanocubes on a mirror (i.e., NCoM, see SI Section S5, Figs. S6, S7). Unlike nanospheres, nanocubes produce more diverse spectral responses, as geometric variations can generate multiple resonances associated with different gap sizes and bottom facet geometries [57, 58, 59]. Their synthesis also tends to yield more complex and less uniform shapes, adding to the spectral diversity. Despite the intrinsically higher variability of NCoM spectra (SI Fig. S6(b)) compared to NPoMs (Fig. 3(d)), SPARX consistently achieved substantially lower prediction errors than a naive statistical baseline (SI Fig. S6(c)), demonstrating that SPARX serves as a general framework capable of handling diverse plasmonic nanostructures beyond spherical geometries.
@@
[…]